OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

2x+y=1 4x+2y=-2 2x+y = 1 2x+y=-1 No solutions